# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ A.3 Gronwall’s Inequality
which completes the proof. Case 2: $\mathcal{N}(\mathbf{u}(0))\leq 0$.
In this case, we may further assume that $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$, since if for some $\overline{t}$, $\mathcal{N}(\mathbf{u}(\overline{t}))>0,$ then we can choose $\overline{t}$ be the new starting time and use the proof for Case 1 to show that $\mathbf{u}(t)$ becomes unbounded at some finite time and also converges in direction. Therefore, we assume $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$. Now, since
@@
we have that $\|\mathbf{u}(t)\|_{2}$ decreases with time, and $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ exists. Now, if $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}=0$, then we have $\lim_{t\rightarrow\infty}\mathbf{u}(t)=\mathbf{0}$ and our proof is complete. Else, suppose $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}=\eta>0$. Then, we have $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq 0$, since $\|\mathbf{u}(t)\|_{2}$ decreases with time. We next show that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$. From Lemma B.4, we know
@@
Thus, $\mathcal{N}(\mathbf{u}(t))$ increases with time. Since we have assumed $\mathcal{N}(\mathbf{u}(t))\leq 0$ for all $t\geq 0$, by monotone convergence we have $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))$ exists. Now, to show $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$, we assume for the sake of contradiction that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=-\gamma<0$. Then, for all $t\geq 0$, we have $\mathcal{N}(\mathbf{u}(t))\geq-\gamma$, since $\mathcal{N}(\mathbf{u}(t))$ is an increasing function of time. Thus,
@@
From the above equation we get that $\|\mathbf{u}(t)\|_{2}$ would become smaller than $\eta$ after some finite amount of time has elapsed. Since $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq 0$, this leads to a contradiction. Therefore, $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$. This also implies $\lim_{t\rightarrow\infty}\tilde{\mathcal{N}}(\mathbf{u}(t))=0$. We next prove that $\lim_{t\rightarrow\infty}\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists. We first define $\hat{\mathbf{u}}(t)=2\mathbf{u}(t)/\eta$. Now, if $\hat{\mathbf{u}}(t)$ converges in direction, then ${\mathbf{u}}(t)$ also converges in direction. To prove $\hat{\mathbf{u}}(t)$ converges in direction we can follow the same approach as in Case 1, specifically from (29) onward. This transformation is essential because, recall, to use Lemma A.1 in Case 1 we had to choose $T$ large enough such that $\|\mathbf{u}(t)\|_{2}>1$, for all $t\geq T$. Here, $\|\mathbf{u}(t)\|_{2}$ may never be greater than $1$, but $\|\hat{\mathbf{u}}(t)\|_{2}\geq 2$, for all $t\geq 0$. So far we have established that either $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges, or $\mathbf{u}(t)$ converges to $\mathbf{0}$. We next show that if $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges to $\mathbf{u}^{*}$, then $\mathbf{u}^{*}$ must be a non-negative KKT point of the constrained NCF. From the proof so far we already know that $\mathcal{N}(\mathbf{u}^{*})=\tilde{\mathcal{N}}(\mathbf{u}^{*})\geq 0.$ Therefore, from Lemma B.3, we only need to show
$L\mathcal{N}(\mathbf{u}^{*})\mathbf{u}^{*}=\nabla\mathcal{N}(\mathbf{u}^{*}),$ (31)
@@
$\left\|\nabla\mathcal{N}(\mathbf{u}^{*})-L\mathcal{N}(\mathbf{u}^{*})\mathbf{u}^{*}\right\|_{2}\geq\gamma.$ (32)
For any $\epsilon>0$, we define the set $\mathbf{u}_{\epsilon}=\{\mathbf{u}:\|\mathbf{u}-\mathbf{u}^{*}\|_{2}\leq\epsilon\}$. Since $\mathcal{N}(\mathbf{u})$ has locally Lipschitz gradient, given $\gamma$, we can choose sufficiently small $\epsilon\in(0,1)$ such that for all $\mathbf{u}\in\mathbf{u}_{\epsilon}$, we have
$\|{\nabla}\mathcal{N}(\mathbf{u})-\nabla\mathcal{N}(\mathbf{u}^{*})\|_{2}\leq\gamma/4.$ (33)
Since $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges to $\mathbf{u}^{*}$ and $\mathcal{N}(\mathbf{u})$ is continuous, we choose $T$ large enough such that for all $t\geq T$,
$\left\|\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}-\mathbf{u}^{*}\right\|_{2}\leq\epsilon,\left\|\frac{L\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)-L\mathbf{u}^{*}\mathcal{N}(\mathbf{u}^{*})\right\|\leq\gamma/4.$ (34)
From the two cases discussed above we know that if $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges, then $\|\mathbf{u}(t)\|_{2}$ either goes to infinity or remains away from 00. Thus, we may also assume that $T$ is large enough such that for some $\eta>0$, we have
$\|\mathbf{u}(t)\|_{2}\geq\eta,\forall t\geq T.$
@@
$\geq$ $\left\|\nabla\mathcal{N}(\mathbf{u}^{*})-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}-\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\nabla\mathcal{N}(\mathbf{u}^{*})\right\|_{2}$
$\geq$ $\left\|\nabla\mathcal{N}(\mathbf{u}^{*})-{L\mathbf{u}^{*}}\mathcal{N}\left({\mathbf{u}^{*}}\right)\right\|_{2}-\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\nabla\mathcal{N}(\mathbf{u}^{*})\right\|_{2}-\left\|{L\mathbf{u}^{*}}\mathcal{N}\left({\mathbf{u}^{*}}\right)-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2},$
where in the second equality we used $(L-1)$-homogeneity of $\nabla\mathcal{N}(\mathbf{u})$ and $L$-homogeneity of $\mathcal{N}(\mathbf{u})$. The inequalities make repeated use of triangle inequality of norms. Hence, using (32), (33) and (34), for all $t\geq T$, we have
@@
Using the above inequality and (25) we have
$\frac{d}{dt}\left(\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)\right)=\left\|\left(\mathbf{I}-\frac{\mathbf{u}(t)\mathbf{u}(t)^{\top}}{\|\mathbf{u}(t)\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u}(t))}{\|\mathbf{u}(t)\|_{2}^{L/2}}\right\|_{2}^{2}\geq\gamma^{2}\|\mathbf{u}(t)\|_{2}^{L-2}/4\geq\gamma^{2}\eta^{L-2}/4.$
However, since the limit of $\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)$ exists, the above inequality leads to a contradiction, which completes our proof. ∎
Before proceeding to proof of Theorem 5.1, we state a useful lemma.
Lemma B.6.
@@
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (36)
where $\mathbf{u}^{*}$ is a non-negative KKT point of
@@
Similar to the proof of Lemma 5.2, we consider two cases. Case 1: $\mathcal{N}(\mathbf{u}(0))>0$.
Let $\mathcal{N}(\mathbf{u}(0))=\gamma>0,$ thus $\|\mathbf{u}(0)\|_{2}>0$. From Lemma B.5, $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(0)\|_{2}$, for all $t\geq 0$, which implies we can choose $\eta=\|\mathbf{u}(0)\|_{2}$ and $T=0$. Next, from Lemma 5.2 and its proof we know that for some finite $T^{*}$, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}^{*}$, where $\mathbf{u}^{*}$ is a non-negative KKT point of the constrained NCF (37). Thus, for any $\epsilon\in(0,1)$, we can choose $T_{\epsilon}\geq T=0$, and large enough but less than $T^{*}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon.$ (38)
@@
In this case, we can also assume that $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$, since if $\mathcal{N}(\mathbf{u}(\overline{t}))>0$, for some $\overline{t}$, then from Lemma B.5, we know $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(\overline{t})\|_{2}$, for all $t\geq\overline{t}$. Moreover, since $\mathcal{N}(\mathbf{u}(\overline{t}))>0$ implies $\|\mathbf{u}(\overline{t})\|_{2}>0$, we can choose $\eta=\|\mathbf{u}(\overline{t})\|_{2}$ and $T=\overline{t}$. Also, in this case, from the proof of Lemma 5.2, we know that for some finite $T^{*}$, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}^{*}$, where $\mathbf{u}^{*}$ is a non-negative KKT point of the constrained NCF (37). Thus, similar to Case 1, for any $\epsilon\in(0,1)$ we can choose $T_{\epsilon}\geq\overline{t}$ and $B_{\epsilon}$ as desired. Hence, suppose $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$. Then, since
@@
Since we have assumed $\lim_{t\rightarrow\infty}\mathbf{u}(t)\neq\mathbf{0},$ we know $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}>0$, and we can choose $\eta=\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ and $T=0$. Further, since $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}^{*}$, where $\mathbf{u}^{*}$ is a non-negative KKT point of the constrained NCF (37), for any $\epsilon\in(0,1)$, we can choose $T_{\epsilon}$ large enough such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}^{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon.$ (39)

### v2

@@ B.1 Proof of Lemma 2
which completes the proof. We next show that there exists $\kappa>0$ such that
$\|\mathbf{u}(t)\|_{2}\geq\frac{\kappa}{(T^{*}-t)^{1/(L-2)}},\forall t\in[0,T^{*}),$ (29)
which will be useful later. Define $\mathbf{h}(t)$ such that
$\mathbf{u}(t)=\frac{\mathbf{h}(t)}{(T^{*}-t)^{1/(L-2)}}.$
Our goal is to show $\|\mathbf{h}(t)\|_{2}\geq\kappa$, for all $t\in[0,T^{*})$ and for some $\kappa>0$. Now, since $\|\mathbf{u}(t)\|_{2}>0$, for all $t\in[0,T^{*})$, therefore, $\|\mathbf{h}(t)\|_{2}>0$, for all $t\in[0,T^{*})$ We next show $\lim_{t\rightarrow T^{*}}\|\mathbf{h}(t)\|_{2}>0.$ Since
$\frac{1}{2}\frac{d\|\mathbf{u}\|_{2}^{2}}{dt}=L\mathcal{N}(\mathbf{u})=L\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\|\mathbf{u}\|_{2}^{L},$
we have
$\frac{1}{\|\mathbf{u}\|_{2}^{L-1}}\frac{d\|\mathbf{u}\|_{2}}{dt}=L\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right).$
Integrating both sides from 00 to $t\in(0,T^{*})$, we get
[…]
Re-arranging the above equation gives us
[…]
Substituting $\mathbf{u}(t)=\mathbf{h}(t)/(T^{*}-t)^{1/(L-2)}$, we get
[…]
In LHS of the above equality, at $t=T^{*}$, the denominator is obviously 00, and the numerator is also 00 since, using eq. 30,
[…]
Therefore, using L’Hopital’s rule,
$\frac{1}{\|\mathbf{h}(T^{*})\|_{2}^{L-2}}=\frac{\lim_{t\rightarrow T^{*}}-L(L-2)\mathcal{N}(\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2})}{-1}=\lim_{t\rightarrow T^{*}}L(L-2)\mathcal{N}(\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2})>0.$
Hence, $\|\mathbf{h}(t)\|_{2}>0$, for all $t\in[0,T^{*}]$, which implies there exists some $\kappa>0$ such that $\|\mathbf{h}(t)\|_{2}\geq\kappa$, for all $t\in[0,T^{*})$ and proving eq. 29.
Case 2: $\mathcal{N}(\mathbf{u}(0))\leq 0$.
In this case, we may further assume that $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$, since if for some $\overline{t}$, $\mathcal{N}(\mathbf{u}(\overline{t}))>0,$ then we can choose $\overline{t}$ be the new starting time and use the proof for Case 1 to show that $\mathbf{u}(t)$ becomes unbounded at some finite time and also converges in direction. Therefore, we assume $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$.
Now, since
@@
we have that $\|\mathbf{u}(t)\|_{2}$ decreases with time, and $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ exists. Now, if $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}=0$, then we have $\lim_{t\rightarrow\infty}\mathbf{u}(t)=\mathbf{0}$ and our proof is complete.
Else, suppose $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}=\eta>0$. Then, we have $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq 0$, since $\|\mathbf{u}(t)\|_{2}$ decreases with time. We next show that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$. From Lemma 12, we know
@@
Thus, $\mathcal{N}(\mathbf{u}(t))$ increases with time. Since we have assumed $\mathcal{N}(\mathbf{u}(t))\leq 0$ for all $t\geq 0$, by monotone convergence we have that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))$ exists. Now, to show $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$, we assume for the sake of contradiction that $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=-\gamma<0$. Then, for all $t\geq 0$, we have $\mathcal{N}(\mathbf{u}(t))\geq-\gamma$, since $\mathcal{N}(\mathbf{u}(t))$ is an increasing function of time. Thus,
@@
From the above equation we get that $\|\mathbf{u}(t)\|_{2}$ would become smaller than $\eta$ after some finite amount of time has elapsed. Since $\|\mathbf{u}(t)\|_{2}\geq\eta$, for all $t\geq 0$, this leads to a contradiction. Therefore, $\lim_{t\rightarrow\infty}\mathcal{N}(\mathbf{u}(t))=0$. This also implies $\lim_{t\rightarrow\infty}\tilde{\mathcal{N}}(\mathbf{u}(t))=0$.
We next prove that $\lim_{t\rightarrow\infty}\mathbf{u}(t)/\|\mathbf{u}(t)\|_{2}$ exists. We first define $\hat{\mathbf{u}}(t)=2\mathbf{u}(t)/\eta$. Now, if $\hat{\mathbf{u}}(t)$ converges in direction, then ${\mathbf{u}}(t)$ also converges in direction. To prove $\hat{\mathbf{u}}(t)$ converges in direction we can follow the same approach as in Case 1, specifically from eq. 27 onward. This transformation is essential because, recall, to use Lemma 6 in Case 1 we had to choose $T$ large enough such that $\|\mathbf{u}(t)\|_{2}>1$, for all $t\geq T$. Here, $\|\mathbf{u}(t)\|_{2}$ may never be greater than $1$, but $\|\hat{\mathbf{u}}(t)\|_{2}\geq 2$, for all $t\geq 0$.
So far we have established that either $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges, or $\mathbf{u}(t)$ converges to $\mathbf{0}$. We next show that if $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges to $\mathbf{u}_{*}$, then $\mathbf{u}_{*}$ must be a non-negative KKT point of the constrained NCF.
From the proof so far we already know that $\mathcal{N}(\mathbf{u}_{*})=\tilde{\mathcal{N}}(\mathbf{u}_{*})\geq 0.$ Therefore, from Lemma 11, we only need to show
$L\mathcal{N}(\mathbf{u}_{*})\mathbf{u}_{*}=\nabla\mathcal{N}(\mathbf{u}_{*}),$ (31)
@@
$\left\|\nabla\mathcal{N}(\mathbf{u}_{*})-L\mathcal{N}(\mathbf{u}_{*})\mathbf{u}_{*}\right\|_{2}\geq\gamma.$ (32)
For any $\epsilon>0$, we define the set $\mathbf{u}_{\epsilon}=\{\mathbf{u}:\|\mathbf{u}-\mathbf{u}_{*}\|_{2}\leq\epsilon\}$. Since $\mathcal{N}(\mathbf{u})$ has locally Lipschitz gradient, given $\gamma$, we can choose sufficiently small $\epsilon\in(0,1)$ such that for all $\mathbf{u}\in\mathbf{u}_{\epsilon}$, we have
$\|{\nabla}\mathcal{N}(\mathbf{u})-\nabla\mathcal{N}(\mathbf{u}_{*})\|_{2}\leq\gamma/4.$ (33)
Since $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges to $\mathbf{u}_{*}$ and $\mathcal{N}(\mathbf{u})$ is continuous, we choose $T$ large enough such that for all $t\geq T$,
$\left\|\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}-\mathbf{u}_{*}\right\|_{2}\leq\epsilon,\left\|\frac{L\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)-L\mathbf{u}_{*}\mathcal{N}(\mathbf{u}_{*})\right\|\leq\gamma/4.$ (34)
From the two cases discussed above we know that if $\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}$ converges, then are two scenarios: $\|\mathbf{u}(t)\|_{2}$ either goes to infinity or $\|\mathbf{u}(t)\|_{2}$ remains finite, bounded away from 00, for all $t\geq 0$ and $\tilde{\mathcal{N}}(\mathbf{u}(t))$ converges to 00. Now, if $\|\mathbf{u}(t)\|_{2}$ goes to infinity, then we may assume $T$ is large enough such that for some $\kappa>0$
$\|\mathbf{u}(t)\|_{2}\geq\frac{\kappa}{(T^{*}-t)^{1/(L-2)}},\forall t\in[T,T^{*}),$ (35)
where $\lim_{t\rightarrow T^{*}}{\mathbf{u}(t)}/{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}_{*}.$ In the second scenario, we may assume that $T$ is large enough such that for some $\eta>0$, we have
$\|\mathbf{u}(t)\|_{2}\geq\eta,\forall t\geq T.$ (36)
@@
$\geq$ $\left\|\nabla\mathcal{N}(\mathbf{u}_{*})-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}-\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\nabla\mathcal{N}(\mathbf{u}_{*})\right\|_{2}$
$\geq$ $\left\|\nabla\mathcal{N}(\mathbf{u}_{*})-{L\mathbf{u}_{*}}\mathcal{N}\left({\mathbf{u}_{*}}\right)\right\|_{2}-\left\|{L\mathbf{u}_{*}}\mathcal{N}\left({\mathbf{u}_{*}}\right)-\frac{L\mathbf{u}}{\|\mathbf{u}\|_{2}}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)\right\|_{2}$
-- $\left\|{\nabla}\mathcal{N}\left(\frac{\mathbf{u}}{\|\mathbf{u}\|_{2}}\right)-\nabla\mathcal{N}(\mathbf{u}_{*})\right\|_{2},$
where in the second equality we used $(L-1)$-homogeneity of $\nabla\mathcal{N}(\mathbf{u})$ and $L$-homogeneity of $\mathcal{N}(\mathbf{u})$. The inequalities make repeated use of triangle inequality of norms. Hence, using eq. 32, eq. 33 and eq. 34, for all $t\geq T$, we have
@@
Using the above inequality and eq. 23 we have
$\frac{d}{dt}\left(\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)\right)=\left\|\left(\mathbf{I}-\frac{\mathbf{u}(t)\mathbf{u}(t)^{\top}}{\|\mathbf{u}(t)\|_{2}^{2}}\right)\frac{{\nabla}\mathcal{N}(\mathbf{u}(t))}{\|\mathbf{u}(t)\|_{2}^{L/2}}\right\|_{2}^{2}\geq\gamma^{2}\|\mathbf{u}(t)\|_{2}^{L-2}/4.$
Now, if $\|\mathbf{u}(t)\|_{2}$ goes to infinity, then, from eq. 35, we have
$\frac{d}{dt}\left(\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)\right)\geq\frac{\gamma^{2}\kappa^{L-2}}{4(T^{*}-t)},\forall t\in[T,T^{*})$
Integrating the above equation from $T$ to $t\in(T,T^{*})$, we get
$\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)-\mathcal{N}\left(\frac{\mathbf{u}(T)}{\|\mathbf{u}(T)\|_{2}}\right)\geq\frac{\gamma^{2}\kappa^{L-2}}{4}\int_{T}^{T^{*}}\frac{1}{(T^{*}-t)}=\frac{\gamma^{2}\kappa^{L-2}}{4}\ln\left(\frac{T^{*}-T}{T^{*}-t}\right).$
As $t$ approaches $T^{*}$, the LHS remains finite but the RHS goes to infinity, leading to a contradiction. In the second scenario, where $\|\mathbf{u}(t)\|_{2}$ remains finite, bounded away from 00, for all $t\geq 0$, we have, from eq. 36,
$\frac{d}{dt}\left(\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)\right)\geq\gamma^{2}\eta^{L-2}/4,\forall t\geq T.$
Integrating the above equation from $T$ to $t\geq T$, we get
$\mathcal{N}\left(\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}\right)-\mathcal{N}\left(\frac{\mathbf{u}(T)}{\|\mathbf{u}(T)\|_{2}}\right)\geq\gamma^{2}\eta^{L-2}(t-T)/4.$
Since $\mathcal{N}\left({\mathbf{u}}/{\|\mathbf{u}\|_{2}}\right)$ is bounded, the above inequality can not be true for sufficiently large $t$, leading to a contradiction. This completes our proof of Lemma 2. ∎
Before proceeding to proof of Theorem 1, we state a useful lemma.
Lemma 14.
@@
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}_{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon,\text{ and }\|\mathbf{u}(t)\|_{2}\leq B_{\epsilon},\forall t\leq T_{\epsilon},$ (38)
where $\mathbf{u}_{*}$ is a non-negative KKT point of
@@
Similar to the proof of Lemma 2, we consider two cases.
Case 1: $\mathcal{N}(\mathbf{u}(0))>0$.
Let $\mathcal{N}(\mathbf{u}(0))=\gamma>0,$ thus $\|\mathbf{u}(0)\|_{2}>0$. From Lemma 13, $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(0)\|_{2}$, for all $t\geq 0$, which implies we can choose $\eta=\|\mathbf{u}(0)\|_{2}$ and $T=0$.
Next, from Lemma 2 and its proof we know that for some finite $T^{*}$, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}_{*}$, where $\mathbf{u}_{*}$ is a non-negative KKT point of the constrained NCF eq. 39. Thus, for any $\epsilon\in(0,1)$, we can choose $T_{\epsilon}\geq T=0$, and large enough but less than $T^{*}$ such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}_{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon.$ (40)
@@
In this case, we can also assume that $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$, since if $\mathcal{N}(\mathbf{u}(\overline{t}))>0$, for some $\overline{t}$, then from Lemma 13, we know $\|\mathbf{u}(t)\|_{2}\geq\|\mathbf{u}(\overline{t})\|_{2}$, for all $t\geq\overline{t}$. Moreover, since $\mathcal{N}(\mathbf{u}(\overline{t}))>0$ implies $\|\mathbf{u}(\overline{t})\|_{2}>0$, we can choose $\eta=\|\mathbf{u}(\overline{t})\|_{2}$ and $T=\overline{t}$. Also, in this case, from the proof of Lemma 2, we know that for some finite $T^{*}$, $\lim_{t\rightarrow T^{*}}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}_{*}$, where $\mathbf{u}_{*}$ is a non-negative KKT point of the constrained NCF eq. 39. Thus, similar to Case 1, for any $\epsilon\in(0,1)$ we can choose $T_{\epsilon}\geq\overline{t}$ and $B_{\epsilon}$ as desired.
Hence, suppose $\mathcal{N}(\mathbf{u}(t))\leq 0,$ for all $t\geq 0$. Then, since
@@
Since we have assumed $\lim_{t\rightarrow\infty}\mathbf{u}(t)\neq\mathbf{0},$ we know $\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}>0$, and we can choose $\eta=\lim_{t\rightarrow\infty}\|\mathbf{u}(t)\|_{2}$ and $T=0$. Further, since $\lim_{t\rightarrow\infty}\frac{\mathbf{u}(t)}{\|\mathbf{u}(t)\|_{2}}=\mathbf{u}_{*}$, where $\mathbf{u}_{*}$ is a non-negative KKT point of the constrained NCF eq. 39, for any $\epsilon\in(0,1)$, we can choose $T_{\epsilon}$ large enough such that
$\frac{\mathbf{u}(T_{\epsilon})^{\top}{\mathbf{u}_{*}}}{\|\mathbf{u}(T_{\epsilon})\|_{2}}\geq 1-\epsilon.$ (41)